Grade 9Mechanics

The speed of a bus increases from 50 kilometre Hawa 270 kilometre Hawa is 5 second find the mean acceleration of the bus

Sangeeta
8 Years agoGrade 9

5 Answers

Arun
8 Years ago
Funal velocity v=70*1000/3600 =19.4 m/s
Initial velocity u =50*1000/3600=13.88 m/s
t= 5 second
acceleration = a = (19.4 -13.88)/5= 5.52/5= 1.104 m/s^2
Adarsh
8 Years ago
DEAR,
For calculating acceleration use 1st equation of motion , v=u+at
here,
v = final velocity =270km/hr
u = initial velocity=50km/hr
t = 5second=5/3600hr
Now, putting the value in equation , 270=50+a.5/3600
or, a=220 x 3600/5=158400km/hr .
REGARDS
Adarsh

DEAR,

For calculating acceleration use 1st equation of motion , v=u+at

here,

v = final velocity =270km/hr

u = initial velocity=50km/hr

t = 5second=5/3600hr

Now, putting the value in equation , 270=50+a.5/3600

or, a=220 x 3600/5=158400km/hr .

REGARDS

Adarsh

Chandan
8 Years ago
mean =220\5
=44
hope it helps please approve if it does
Gitanjali Rout
8 Years ago
Funal velocity v=70*1000/3600 =19.4 m/s
Initial velocity u =50*1000/3600=13.88 m/s
t= 5 second
acceleration = a = (19.4 -13.88)/5= 5.52/5= 1.104 m/s^2/////////////////////kjkhiu    jgj
Gitanjali Rout
8 Years ago
DEAR,For calculating acceleration use 1st equation of motion , v=u+athere,v = final velocity =270km/hru = initial velocity=50km/hrt = 5second=5/3600hrNow, putting the value in equation , 270=50+a.5/3600or, a=220 x 3600/5=158400km/hr .